OpenStudy (anonymous):

Find the length and width of the rectangle whose vertices are given. V(-1, -2), W(3, -6), X(9, 0), Y(5, 4) vy= sq.rt. vw= sq.rt.

OpenStudy (anonymous):

VY = \[\sqrt{(5+1)^2 + (4+2)^2}= \sqrt{(6)^2 + (8)^2} = \sqrt{36+ 64} = \sqrt{100} = 10\] VW = \[\sqrt{(3+ 1)^2 + (-6+2)^2} = \sqrt{(4)^2 + (-4 )^2 } = \sqrt{16 + 16} = \sqrt{32}\]